1/1

Manfred Hausmann,  Manfred Hausmann,
Vorfrühling  Vorfrühling

translated by Manfred Retzlaff

Hinter einer halbverfallnen Mauer blüht ein Pflaumenbaum trotz nächtlichem Frost. Es ist von ferne zu sehen, dass es kein Schnee ist, denn süßer Blumenduft weht heran.

Argiope information: The view of this version is restricted. Presumably there are copyright issues.

Translation of the German poem “Vorfrühling” by Manfred Hausmann (∗1898-09-10 – †1986-08-06) into German by Manfred Retzlaff (∗1938-11-04).

Arg-901-1878 (2013-11-22 18:33:36)

Wörtiche Rückübersetzung der Esperanto-Fassung.

Author of this German poem is Manfred Hausmann (∗1898-09-10 – †1986-08-06).

Arg-901-1797 (2013-08-06 23:14:15)

Verkita laŭ poemo de la ĉina poeto Wang An-Schi, Wáng Ānshí. Siehe auch: http: // de. wikipedia. org/ wiki/ Wang_Anshi . Vidu ankaŭ: http: // www. mountainsongs. net/ poet_.php? id= 213 .

http://www.poezio.net/version?poem-id=901&version-id=1878,1797